OpenStudy (dugalde):

Hello can you help me find the complex cube roots of 8(cos(4pi/5)+isin(4pi/5))

zepdrix (zepdrix):

\[\Large\rm z=8\left[\cos\left(\frac{4\pi}{5}\right)+\mathcal i \sin\left(\frac{4\pi}{5}\right)\right]\]\[\Large\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{5}\right)+\mathcal i \sin\left(\frac{4\pi}{5}\right)\right]^{1/3}\]

zepdrix (zepdrix):

Hey There Diana :) Do you remember De'Moivre's Theorem?

zepdrix (zepdrix):

Alright, I'll leave some notes in case you come back. Before we apply De'Moivre's Theorem, let's rewrite our angles inside of our sine and cosine, we need to allow for rotation.\[\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{5}+2k \pi\right)+\mathcal i \sin\left(\frac{4\pi}{5}2k \pi\right)\right]^{1/3}, \quad k=0,\pm1,\pm2,\pm3,...\]We can add any number of full spins. We're allowing our angle to be 4pi/5 or 14pi/5 or 24pi/5 and so on.... because when we take the sine and cosine of those angles, they give us the same values that 4pi/5 would give. Since this is the `third root`, we only want the first three positive values for k, starting from 0, after that our solutions will start to repeat.

zepdrix (zepdrix):

\[\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{5}+2k \pi\right)+\mathcal i \sin\left(\frac{4\pi}{5}2k \pi\right)\right]^{1/3}, \quad k=0,1,2\]Applying De'Moivre's Theorem lets us bring the exponent in as a factor on the angle.\[\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{3\cdot5}+\frac{2k \pi}{3}\right)+\mathcal i \sin\left(\frac{4\pi}{3\cdot5}+\frac{2k \pi}{3}\right)\right], \quad k=0,1,2\]

zepdrix (zepdrix):

Simplify, then plug in your k's to get the three results :)

zepdrix (zepdrix):

I'll show you the first one just so it's clear. Err I'll show you k=2 actually.

zepdrix (zepdrix):

\[\rm z_2^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{3\cdot5}+\frac{2\cdot2 \pi}{3}\right)+\mathcal i \sin\left(\frac{4\pi}{3\cdot5}+\frac{2\cdot2 \pi}{3}\right)\right]\]\[\rm z_2^{1/3}=2\left[\cos\left(\frac{4\pi}{15}+\frac{20\pi}{15}\right)+\mathcal i \sin\left(\frac{4\pi}{15}+\frac{20\pi}{15}\right)\right]\]\[\rm z_2^{1/3}=2\left[\cos\left(\frac{24\pi}{15}\right)+\mathcal i \sin\left(\frac{24\pi}{15}\right)\right]\]\[\rm z_2^{1/3}=2\left[\cos\left(\frac{8\pi}{5}\right)+\mathcal i \sin\left(\frac{8\pi}{5}\right)\right]\]Mmm that doesn't work out so nice -_- hmm probably need a calculator from here.\[\Large\rm \rm z_2^{1/3}=2\left[0.309-\mathcal i 0.951\right]\]\[\Large\rm \rm z_2^{1/3}=0.6180-1.902\mathcal i\]

zepdrix (zepdrix):

So that would be one of our solutions :U

OpenStudy (dugalde):

Thank you so much!

zepdrix (zepdrix):

Did you find the other roots bunny? c:

OpenStudy (dugalde):

not really

zepdrix (zepdrix):

the process too confusing? :o

OpenStudy (anonymous):

ha

OpenStudy (dugalde):

yeah

zepdrix (zepdrix):

hmm :d

OpenStudy (dugalde):

@zepdrix Can you help me find the other two please.

zepdrix (zepdrix):

When k=0, it's pretty straight forward. It's the equation you started with,\[\Large\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{5}\right)+\mathcal i \sin\left(\frac{4\pi}{5}\right)\right]^{\color{orangered}{1/3}}\]no rotation. And apply De'Moivre's Theorem,\[\Large\rm z^{1/3}=8^{1/3}\left[\cos\left(\color{orangered}{\frac{1}{3}}\cdot\frac{4\pi}{5}\right)+\mathcal i \sin\left(\color{orangered}{\frac{1}{3}}\cdot\frac{4\pi}{5}\right)\right]\]And let your calculator do the rest :o

zepdrix (zepdrix):

When k=1, we'll add 1 full spin to our angle.\[\Large\rm z^{1/3}=8^{1/3}\left[\cos\left(\frac{4\pi}{5}+2\pi\right)+\mathcal i \sin\left(\frac{4\pi}{5}+2\pi\right)\right]^{\color{orangered}{1/3}}\]Applying De'Moivre,\[\large\rm z^{1/3}=8^{1/3}\left[\cos\left(\color{orangered}{\frac{1}{3}\cdot}\frac{4\pi}{5}+\color{orangered}{\frac{1}{3}\cdot}2\pi\right)+\mathcal i \sin\left(\color{orangered}{\frac{1}{3}\cdot}\frac{4\pi}{5}+\color{orangered}{\frac{1}{3}\cdot}2\pi\right)\right]\]Simplify, and calculator :O

OpenStudy (dugalde):

how would I put that in my calculator

OpenStudy (dugalde):

@zepdrix I think I got it k 0=2.0004+0.0292i and k 1=1.9972+0.102i